Atsie (atsie):

Some of us know that conditional statements are basically written in "if - then - form." So with that being said, could someone give me a STEP BY STEP (I'm emphasizing because most of the time people just give me the answer and expect a medal) explanation on the problem below? x = 8, x2 - 3x = 40. I'd so appreciate it!! :)

OpenStudy (anonymous):

Ok first you want to replace x so 8*3-3*40 did u get that?

OpenStudy (anonymous):

Then you want to solve it so, 8*3=24 3*40=120 120-24=96

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

@Astie

OpenStudy (anonymous):

@Atsie sorry

Atsie (atsie):

Yes it does in fact! I have sort of an idea with where I am going I guess. So these are the options I am given to answer the above question in conditional statement form. A. If x = 8, then x2 - 3x = 40 B. If x2 - 3x = 40, then x = 8. C. If x2 - 3x = 40, then x(x - 3) = 40 D. If x(x - 3) = 40, then x2 - 3x = 40 What do you personally think is the best one to pick and why? I'm trying to make sure I understand things correctly.

Atsie (atsie):

@sydneymmelin :)

OpenStudy (anonymous):

I would have to go with A because when u are answering a problem like this u need to make sure you first go what x is then you have your equation. Is that helpful. @Atsie

Atsie (atsie):

Indeed it is. Thank you lovely! Medal is all yours :)

OpenStudy (anonymous):

LOL. Thnx.

Atsie (atsie):

No problem :))